# Radio-frequency cascade readout and coherent exchange control of spin qubits fabricated using a 300 mm wafer process

### v3

@@ Abstract
Leveraging advanced semiconductor manufacturing promises significant potential for scaling up silicon-based quantum processors. In this context, control and readout of individual spin qubits have been shown on devices manufactured on 300 mm wafer metal-oxide-semiconductor (MOS) processes, yet quantum processors require two-qubit interactions to operate. Here, we use a 300 mm natural silicon MOS process customized for spin qubits and demonstrate coherent control of two electron spins using the spin-spin exchange interaction, which forms the basis for entangling gates such as $\sqrt{\text{SWAP}}$. We observe gate dephasing times of up to $T_{2}^{*}\approx 500$ ns and a gate quality factor of 10. For readout, we introduce a novel dispersive readout technique, the radio-frequency electron cascade, that simplifies the qubits unit cell while providing high gain. We achieve a 74% singlet-triplet readout fidelity in 8 $\upmu$s integration time, the highest-performing dispersive readout demonstration on a planar MOS process. The combination of sensitive dispersive readout with industrial-grade manufacturing marks a crucial step towards the large-scale integration of silicon quantum processors.
@@
Spin qubits based on silicon metal-oxide-semiconductor (MOS) technology  [20, 21] and Si/SiGe heterostructures [22, 23, 16, 11, 24] are of particular interest for industrial manufacturing. Specifically, the MOS approach shares similarities with modern silicon field-effect transistor (FET) manufacturing, enabling the formation of MOS QDs in both planar devices [13, 20, 21, 25, 26, 27], and etched silicon structures, such as nanowires [28, 29, 30] or finFETs [6, 31]. While the latter constrains the qubit topology to 2xN bilinear QD arrays, the former offers easier scalability towards two-dimensional QD arrays [32], which are essential for implementing quantum error correcting codes such as the surface code. This potential for scaling has motivated the translation of the well-established operation of high-quality qubits within MOS academic devices [8, 13, 25, 20, 26], to demonstrations of single-qubit performance in MOS devices fabricated using semiconductor manufacturing lines [28, 31]. Yet, quantum processors require two-qubit interactions to operate.
For further scale-up, methods that simplify the readout infrastructure must be developed. The current standard for sensing, the radio-frequency single-electron transistor, offers high-fidelity readout [3, 5] at the cost of occupying substantial space on the qubit chip, which limits qubit connectivity. To address this limitation, efforts have focused on integrating fast and compact radio-frequency measurement techniques [33, 34, 35], which substantially reduce the readout footprint. However, despite recent advances in other areas, dispersive readout metrics have seen little progress in planar MOS in recent years [34].
@@
In this Article, we study electron spins within a silicon double quantum dot (DQD) device fabricated using a 300 mm wafer line [36, 37] to address the two aforementioned outstanding challenges: First, we introduce a new in situ dispersive sensing mechanism (termed radiof-requency- electron cascade) which offers enhanced sensitivity for spin qubit readout. We achieve a minimum integration time of 13 $\upmu$s and using a physical model, we calculate 74% singlet-triplet readout fidelity in an 8 $\upmu$s integration time. Second, we demonstrate control of an exchange-mediated coherent interaction, which forms the basis for a $\sqrt{\rm SWAP}$ gate between two spin qubits.
Radio-frequency electron cascade readout
@@
Reading out spin qubits within semiconductor QDs typically involves mapping a spin state of interest onto a charge state of one or more QDs [38, 39], which can then be detected using a variety of charge sensing methods [40]. For example, Pauli spin blockade (PSB) can be used to map the singlet and triplet states of a pair of spin qubits onto two different charge configurations of a DQD (e.g. (1,1) or (0,2)), which are then detected by a single electron transistor [41] or single electron box [4]. In-situ dispersive readout of a DQD combines these into a single step, using PSB to directly distinguish between singlet and triplet states through their difference in AC polarisability [42]. This difference in polarisability is detected by incorporating the DQD into a radio-frequency (rf) tank circuit and measuring changes in the reflected rf signal. However, in-situ dispersive readout has suffered from low sensitivity in planar MOS silicon quantum devices due to the relatively low gate lever arms [34]. To improve the sensitivity of this technique, we introduce a third quantum dot (QD) coupled to a charge reservoir, which acts as an amplifier in measuring the AC-polarisability of the DQD. Instead of measuring the usual single-electron alternating current generated by the cyclic tunneling between the two-spin singlet states of the DQD [40], we leverage the synchronised single-electron AC current at the third dot-reservoir system generated as a consequence of the strong capacitive coupling to the DQD. Our approach offers the benefit of charge enhancement techniques such as latching [43], DC cascading [44] and spin-polarized single-electron boxes [45] while retaining the non-demolition nature of in-situ dispersive readout methods [46].
@@
Fig. 2: |||| Radio-frequency cascade readout fidelity. a Charge stability diagram of the DQD around the (1,1)-(0,2) interdot charge transition (ICT), with a voltage pulse sequence and detuning axis $\varepsilon$ overlaid b, Reflected resonator voltage $V_{\rm rf}$ as a function of detuning across the ICT around $\varepsilon=0$, fit to eq. (5) (dashed orange line), using tunnel coupling $t_{c}$. c Energy diagram showing dependence of the two-electron spin states, singlet |$\ket{\text{S}}$ and triplet |$\ket{\text{T}}$ as a function of $\varepsilon$. The pulse sequence depicts the initialisation (I) of the |$\ket{{\rm S(0,2)}}$ via an adiabatic ramp from the $(0,1)$ empty (E) state, followed by a non-adiabatic pulse to $\varepsilon=0$ for measurement (M). d, $V_{\rm rf}$ as a function of wait time $\tau_{\rm M}$ at $\varepsilon=0$ before measurement, following the pulse sequence depicted in panels a and c. The fitted line indicates a |$\ket{\text{S}}$ to |$\ket{{\rm T_{-}(1,1)}}$ relaxation time $T_{1}=24~{}\upmu$s. e, Calculated readout infidelity $1-\mathcal{F}_{r}$ as a function of integration time $\tau_{\text{int}}$ as described by Eq. (2). Solid lines depict relaxation time $T_{1}$ limited fidelity, whereas dashed lines depict $T_{1}=\infty$. Parameters used from left to right include: $\tau_{\text{min}}=\left\{94~{}\text{ns},13~{}\upmu\text{s},"\right\}$, $T_{1}=\left\{24~{}\upmu\text{s},",4.5~{}\text{ms}\right\}$.
@@
where $\alpha_{{\rm R},j}$ represents the gate lever arm between the reservoir and QD $j$. In particular, the sensor detects not only the interdot gate polarization charge $(\alpha_{\text{R,1}}-\alpha_{\text{R,2}})e$ but also the cascaded charge collected at the reservoir, $(1-\alpha_{\text{R,ME}})e$, (see Supplementary Section S2 for derivation). By comparing the measured signal-to-noise ratio (SNR) with and without the cascade effect, we find a lower bound for the power amplification factor, $A\geq 1010\pm 40$ $(+36~{}\text{\rm dB})$. The cascade SNR is extracted from the interdot charge transition fit shown in Fig. 2(a-b) (see eq. (5-6) in Methods). For the case where the cascade is absent, the SNR is assumed to be 0.5, as there is no observable signal, as shown in Supplementary Fig. S1(a) and (e). From the cascade SNR we extract a minimum integration time of $\tau_{\rm min}=13.0\pm 0.5~{}\upmu{\rm s}$ (see eq. (7) in Methods), representing an improvement of over two orders-of-magnitude on prior planar MOS demonstrations of in-situ dispersive readout [34].
@@
We use the rf-cascade to distinguish between the singlet and triplet states of the DQD via PSB. The signature of PSB can be observed by measuring the asymmetric disappearance of the interdot charge transitions as a function of increasing the applied magnetic field [42], as shown in the inset of Fig. 1(c). At low magnetic field $(B\leq 200~{}\text{\rm mT})$, the system is free to oscillate between singlet states (|$\ket{\text{S}(1,1)}\leftrightarrow\ket{\text{S}(0,2)}$) due to the action of the rf drive, yielding a signal in the rf response. However, at higher fields $(B\geq 200~{}\text{\rm mT})$, the polarised triplet |$\ket{\text{T}_{-}(1,1)}$ state becomes the ground state for $\varepsilon\geq 0$ (as shown in Fig. 2(c)), preventing a charge transition and resulting in the disappearance of the signal, initially for the region of the transition closer to the $(1,1)$ charge configuration. A quantum capacitance-based simulation of the data in the inset suggests an inter-dot tunnel coupling of $t_{c}=2.4$ GHz and electron temperature $T_{e}=50$ mK (see Supplementary Section S3). Figure 2(d) shows the decay from |$\ket{\text{S}}$ to |$\ket{\text{T}_{-}(1,1)}$ at $\varepsilon=0$ for applied magnetic field $B=250~{}$mT. The corresponding relaxation time is $T_{1}=24\pm 7~{}\upmu$s, which is consistent with a study of a similar device, where a relaxation time of $T_{1}=10~{}\upmu{\rm s}$ was observed near zero detuning for a comparable tunnel coupling ($t_{c}\geq 1.9~{}$GHz) [51]. In that study, $T_{1}$ was found to vary exponentially with $t_{c}$, with relaxation times exceeding $200~{}\upmu$s for $t_{c}\leq 1~{}$GHz.
We assess the performance of the radio-frequency driven electron cascade by calculating the readout fidelity, $\mathcal{F}_{r}$ as a function of integration time, using the expression [40, 52],
$\mathcal{F}_{r}=\frac{1}{2}\left[1+\text{erf}\left(\sqrt{\frac{\tau_{\text{int}}}{2\tau_{\text{min}}}}\right)\exp\left(\frac{\tau_{\text{int}}}{2T_{1}}\right)\right].$ (2)
@@
We calculate a readout fidelity of $\mathcal{F}_{r}=74\begin{subarray}{c}+1\\ -2\end{subarray}\%$ based on the experimentally obtained $\tau_{\rm min}=13~{}\upmu$s and integration time $\tau_{\text{int}}=8~{}\upmu$s. The corresponding readout infidelity $1-\mathcal{F}_{r}$ is depicted in Fig. 2(c) (dark purple line). This fidelity is comparable to prior in-situ readout demonstrations, but is achieved with integration times that are 1-2 orders of magnitude faster than those reported in previous silicon planar MOS and implanted donor systems [34, 33]. Our performance approaches that of integrated on-chip resonators in Si/SiGe [35], but without the added fabrication complexity of on-chip resonator integration. To achieve further improvements in fidelity, two strategies are indicated in Fig. 2(c): ($i$) reducing the minimum integration time to 949494~{}94ns, to achieve a fidelity of $99\%$ (light purple line), by enhancing the SNR through resonator optimisation or in-situ matching techniques [40, 53]; and ($ii$) by extending the relaxation time of the system [51] to that reported in the state-of-the-art MOS [34] (orange line), resulting in an estimated fidelity of $99.2\%$.
@@
We highlight that, unlike other charge enhancement techniques [43, 45, 44], the rf electron cascade retains the non-demolition nature of in-situ dispersive readout measurements since the DQD system remains in an eigenstate after a measurement is performed [54, 1]. We note that the rf excitation is continuously applied for all measurements in this Article.
@@
Having established a method to distinguish between singlet and triplet spin states, our goal is to prepare and coherently control spin states of the DQD through voltage pulses along the detuning axis, $\epsilon$. Such pulses bring the DQD: i) from the (0,2) charge configuration in which a singlet is prepared; ii) into the (1,1) region where the electron spins are spatially separated between QDs and may evolve; and iii) back to an intermediate point where they can be measured (see Fig. 3(a-c)). Deep in the (1,1) region, the spin basis states are predominantly $\ket{\uparrow\downarrow}$ and $\ket{\downarrow\uparrow}$. Under an adiabatic ramp to $\epsilon=0$ for readout, these two states map onto the |$\ket{\text{T}_{0}(1,1)}$ and |$\ket{\text{S}(0,2)}$, respectively. The basis states are separated in energy by $h\Omega=\sqrt{J(\varepsilon)^{2}+\Delta E_{z}^{2}}$, where we include the kinetic exchange interaction $J(\varepsilon)$ and the Zeeman energy difference between electrons in each dot $\Delta E_{z}$. The spin detuning $\Delta E_{z}=\Delta g\mu_{B}B+g\mu_{B}\Delta B_{\text{HF}}$ (where $\mu_{B}$ is the Bohr magneton, $h$ is Planck’s constant) contains two main contributions: (i) the difference in g-factor between QDs $\Delta g=\left|g_{2}-g_{1}\right|$ arising from variations in the spin-orbit interaction (SOI) present near the Si/SiO2 interface  [55, 56, 26]; and (ii) the difference in the effective 29Si nuclear magnetic field experienced by each QD, $\Delta B_{\text{HF}}$. The random fluctuations in the effective magnetic field experienced by each electron in the DQD can be described by a normal distribution with mean of 0 (given the negligible spin polarisation) and standard deviation $\sigma_{\text{HF}}=30\pm 4~{}\upmu$T, as we shall see later. This value corresponds to a hyperfine energy strength of $3.4\pm 0.4$ neV, which aligns well with other reports in natural silicon [22, 48, 57].
@@
The entangling two-qubit gate achieved between the spin qubits under the exchange interaction depends on the ratio $J/\Delta E_{z}$, tending to a $\sqrt{\text{SWAP}}$ operation as $J\gg\Delta E_{z}$ or a C-PHASE when $J\ll\Delta E_{z}$, though any gate within this set parameterised by $J/\Delta E_{z}$ can be used as the building block for a quantum error correcting code such as the surface code [63]. Defining the qubit quality factor as $Q=T_{2}^{*}\Omega$ (i.e. the number of periods before the amplitude of oscillations decays by $1/e$), we find $Q\gtrapprox 10$ in the region $J/\Delta E_{z}=2.1-3.2$, on par with prior reports across a range of semiconductor platforms [26, 47, 48, 58, 64, 65]. This provides an upper bound estimate on the achievable two-qubit gate fidelity using the approximation $\mathcal{F}\approx 1-1/4Q\lessapprox 98\%$ [66]. To implement error-correctable two-qubit gates this fidelity would need to surpass $99\%$ [15], which could be achieved using isotopically enriched silicon. In the next section, we extend the coherence time using spin refocusing techniques.
@@
In the exchange echo sequence, after initialising a |$\ket{\text{S}(1,1)}$ state and applying a $X^{\prime}_{\pi/2}$ rotation, the two-electron system dephases under the effect of charge noise for a time $\tau_{1}$. The free evolution occurs at a detuning point where $J/\Delta E_{z}=12.8\pm 1.6$ where we measured $T_{2}^{*}=43\pm 3~{}$ns (see Fig 4(f)). We then refocus the spins by applying a $X^{\prime}_{\pi}$ rotation and let the system evolve for $\tau_{2}$ until a second $X^{\prime}_{\pi/2}$ rotation maps the resulting state to the |$\ket{\text{S}}$-|$\ket{\text{T}_{0}}$ axis. We extract the amplitude of the echo by fitting the signal in the $\tau_{2}-\tau_{1}$ domain (see Fig. 5(c)), and plot its value as a function of total free evolution time $\tau_{1}+\tau_{2}$. Fig. 5(d) shows the echo amplitude decays exponentially with total time ($\tau_{2}+\tau_{1}$), yielding a characteristic $T_{2,J}^{\text{echo}}=0.42\pm 0.02~{}\upmu$s which corresponds to an order of magnitude increase in the coherence time, similar to prior work with silicon devices [67, 26, 64]. From our fit to Eq. (4), we extract a magnetic-noise-limited $T_{2,\Delta E_{z}}^{*}=3.3~{}\upmu$s at this set point, indicating that sources other than magnetic noise limit $T_{2,J}^{\text{echo}}$. We speculate that residual high frequency charge noise limits the echo coherence [7], and may include factors such as the effect of the rf tone which is on throughout the control sequence.

### v4

# Radio-frequency cascade readout of coupled spin qubits fabricated using a 300 mm wafer process

@@ Abstract
Advanced semiconductor manufacturing offers a promising path to scaling up silicon-based quantum processors by improving yield, uniformity, and integration. Individual spin qubit control and readout have been demonstrated in quantum dots fabricated using 300 mm wafer metal-oxide-semiconductor (MOS) processes, yet quantum processors require two-qubit interactions to operate. Here, we use a 300 mm natural silicon MOS process customised for spin qubits and demonstrate coherent control of two electron spins using the exchange interaction, forming the basis for entangling gates such as $\sqrt{\text{SWAP}}$. We measure gate dephasing times of up to $T_{2}^{*}\approx 500$ ns with a quality factor of 10. For readout, we introduce a novel dispersive readout technique, the radio-frequency electron cascade, that simplifies the qubit unit cell while providing high gain. This method achieves an ${\rm SNR}$ of 6 in 46 $\upmu$s, the highest-performing dispersive readout demonstration in a planar MOS process. The combination of sensitive dispersive readout with industrial-grade manufacturing marks a crucial step towards large-scale integration of silicon quantum processors.
@@
Spin qubits based on silicon metal-oxide-semiconductor (MOS) technology  [20, 21] and Si/SiGe heterostructures [22, 23, 16, 11, 24] are of particular interest for industrial manufacturing. Specifically, the MOS approach shares similarities with modern silicon field-effect transistor (FET) manufacturing, enabling the formation of MOS QDs in both planar devices [13, 20, 21, 25, 26, 27], and etched silicon structures, such as nanowires [28, 29, 4] or finFETs [6, 30]. While the latter constrains the qubit topology to 2xN bilinear QD arrays, the former offers easier scalability towards two-dimensional QD arrays [31], which are essential for implementing quantum error correcting codes such as the surface code [15]. This potential for scaling has driven the translation of the well-established operation of high-quality qubits within academic MOS devices [8, 13, 25, 20, 26], to demonstrations of single-qubit performance in MOS devices fabricated using semiconductor manufacturing lines [28, 30]. Yet, quantum processors require two-qubit interactions to operate.
To enable further scaling, methods that simplify the readout infrastructure must be developed. The current standard for sensing, the radio-frequency single-electron transistor, offers high-fidelity readout [3, 5] at the cost of occupying substantial space on the qubit chip, which limits qubit connectivity. To address this limitation, efforts have focused on integrating fast and compact radio-frequency measurement techniques [32, 33, 34], which substantially reduce the readout footprint. However, despite recent advances in other areas, dispersive readout metrics have seen little progress in planar MOS in recent years [33].
@@
In this Article, we study electron spins within a silicon double quantum dot (DQD) device fabricated using a 300 mm wafer line [35, 36] to address the two aforementioned outstanding challenges: First, we introduce a new in situ dispersive sensing mechanism (termed radio-frequency electron cascade) which offers enhanced sensitivity for spin qubit readout. We achieve a minimum integration time of 8 $\upmu$s and, using a physical model [37], we calculate a 67% singlet-triplet readout fidelity limited by spin relaxation. Second, we demonstrate control of an exchange-mediated coherent interaction, which forms the basis for a $\sqrt{\rm SWAP}$ gate between two spin qubits.
Radiofrequency electron cascade readout
@@
Reading out spin qubits within semiconductor QDs typically involves mapping a spin state of interest onto a charge state of one or more QDs [38, 39], which can then be detected using a variety of charge sensing methods [40]. For example, Pauli spin blockade (PSB) can be used to map the singlet and triplet states of a pair of spin qubits onto two different charge configurations of a DQD (e.g. (1,1) or (0,2)), which are then detected by a single electron transistor [41] or single electron box [4]. In-situ dispersive readout of a DQD combines these into a single step, using PSB to directly distinguish between singlet and triplet states through their difference in AC polarisability [42]. This difference in polarisability is detected by incorporating the DQD into a radiofrequency (rf) tank circuit and measuring changes in the reflected rf signal. However, in-situ dispersive readout has suffered from low sensitivity in planar MOS silicon quantum devices due to the relatively low gate lever arms [33]. To improve the sensitivity of this technique, we introduce a third quantum dot (QD) coupled to a charge reservoir, which acts as an amplifier in measuring the AC-polarisability of the DQD. Instead of measuring the usual single-electron alternating current generated by the cyclic tunneling between the two-spin singlet states of the DQD [40], we leverage the synchronised single-electron AC current at the third dot-reservoir system generated as a consequence of the strong capacitive coupling to the DQD. Our approach offers the benefit of charge enhancement techniques such as latching [43], DC cascading [44] and spin-polarized single-electron boxes [45] while retaining the non-demolition nature of in-situ dispersive readout methods [46].
@@
Fig. 2: |||| Radio-frequency cascade readout fidelity. a Charge stability diagram of the DQD around the (1,1)-(0,2) interdot charge transition (ICT), with a voltage pulse sequence and detuning axis $\varepsilon$ overlaid b Energy diagram showing dependence of the two-electron spin states, singlet |$\ket{\text{S}}$ and triplet |$\ket{\text{T}}$ as a function of $\varepsilon$. The pulse sequence depicts the initialisation (I) of the |$\ket{{\rm S(0,2)}}$ via an adiabatic ramp from the $(0,1)$ empty (E) state, followed by a non-adiabatic pulse to $\varepsilon=0$ for measurement (M). c, Reflected resonator voltage $V_{\rm rf}$ as a function of detuning across the ICT around $\varepsilon=0$, fit to Eq. (4) (dashed orange line). d, Magneto-spectroscopy of the interdot charge transition of the DQD along detuning, $\varepsilon$. e, $V_{\rm rf}$ as a function of wait time $\tau_{\rm M}$ at $\varepsilon=0$ before measurement, following the pulse sequence depicted in panels a and c. The fitted line indicates a |$\ket{\text{S}}$ to |$\ket{{\rm T_{-}(1,1)}}$ relaxation time $T_{1}=24~{}\upmu$s. f, Calculated readout infidelity $1-\mathcal{F}_{r}$ as a function of integration time $\tau_{\text{int}}$ as described by Eq. (7). Solid lines depict relaxation time $T_{1}$ limited fidelity, whereas dashed lines depict $T_{1}=\infty$. Parameters used from left to right include: $\tau_{\text{min}}=\left\{23~{}\text{ns},8~{}\upmu\text{s},"\right\}$, $T_{1}=\left\{24~{}\upmu\text{s},",4.5~{}\text{ms}\right\}$.
@@
where $\alpha_{{\rm R},j}$ represents the gate lever arm between the reservoir and QD $j$. In particular, the sensor detects not only the interdot gate polarization charge $(\alpha_{\text{R,1}}-\alpha_{\text{R,2}})e$ but also the cascaded charge collected at the reservoir, $(1-\alpha_{\text{R,ME}})e$, (see Supplementary Section S2 for derivation). By comparing the measured signal-to-noise ratio (SNR) with and without the cascade effect, we find a lower bound for the power amplification factor, $A\geq(3.4\pm 0.1)\times 10^{3}$ $(+35.4~{}\text{\rm dB})$. The cascade SNR is extracted from the interdot charge transition fit shown in Fig. 2(a-b) (see Eq. (4-5) in Methods). For the case where the cascade is absent, the SNR is assumed to be 0.5, as there is no observable signal, as shown in Supplementary Fig. S1(a) and (e). From the cascade SNR we extract a minimum integration time of $\tau_{\rm min}=7.6\pm 0.2~{}\upmu{\rm s}$ (see Eq. (6) in Methods), representing an improvement of over two orders-of-magnitude on prior planar MOS demonstrations of in-situ dispersive readout [33].
@@
We use the rf-cascade to distinguish between the singlet and triplet states of the DQD via PSB. The signature of PSB can be observed by measuring the asymmetric disappearance of the interdot charge transitions as a function of increasing the applied magnetic field [42], as shown Fig. 2(c). At low magnetic field $(B\leq 200~{}\text{\rm mT})$, the system is free to oscillate between singlet states (|$\ket{\text{S}(1,1)}\leftrightarrow\ket{\text{S}(0,2)}$) due to the action of the rf drive, yielding a signal in the rf response. However, at higher fields $(B\geq 200~{}\text{\rm mT})$, the polarised triplet |$\ket{\text{T}_{-}(1,1)}$ state becomes the ground state for $\varepsilon\geq 0$ (as shown in Fig. 2(d)), preventing a charge transition and resulting in the disappearance of the signal, initially for the region of the transition closer to the $(1,1)$ charge configuration. A quantum capacitance-based simulation of the data in the inset suggests an inter-dot tunnel coupling of $t_{c}=2.4$ GHz and electron temperature $T_{e}=50$ mK (see Supplementary Section S4). Figure 2(e) shows the decay from |$\ket{\text{S}}$ to |$\ket{\text{T}_{-}(1,1)}$ at $\varepsilon=0$ for applied magnetic field $B=250~{}$mT. The corresponding relaxation time is $T_{1}=24\pm 7~{}\upmu$s, which is consistent with a study of a similar device, where a relaxation time of $T_{1}=10~{}\upmu{\rm s}$ was observed near zero detuning for a comparable tunnel coupling ($t_{c}\geq 1.9~{}$GHz) [51]. In that study, $T_{1}$ was found to vary exponentially with $t_{c}$, with relaxation times exceeding $200~{}\upmu$s for $t_{c}\leq 1~{}$GHz. Lowering $t_{c}$ would also enhance the readout sensitivity, as the dispersive signal is maximised at $22t_{c}=f_{\text{rf}}/$ 2 [peri2024polarimetryspinssolidstate], where $f_{\text{rf}}$ is the resonator drive frequency, which we set to 512 MHz.
@@
We assess the performance of the radio-frequency driven electron cascade by calculating the readout fidelity, $\mathcal{F}_{r}$ as a function of integration time, using Eq. (7) in Methods [37]. We calculate a relaxation-limited readout fidelity of $\mathcal{F}_{r}=67\pm 1\%$ based on the experimentally obtained $\tau_{\rm min}=7.6~{}\upmu$s and integration time $\tau_{\text{int}}=8~{}\upmu$s. The corresponding readout infidelity $1-\mathcal{F}_{r}$ is depicted in Fig. 2(f) (dark purple line). This infidelity is a comparable order of magnitude to prior in-situ readout demonstrations, but is achieved with integration times that are 1-2 orders of magnitude faster than those reported in previous silicon planar MOS and implanted donor systems [33, 32] (see Supplementary Table 1). To achieve a $99\%$ fidelity, two strategies are indicated in Fig. 2(f): ($i$) reducing the minimum integration time to 242424~{}24ns (light purple line), by enhancing the SNR through resonator optimisation or quantum-limited amplification [40, 52]; and ($ii$) by extending the relaxation time of the system, which for the $T_{1}=4.5~{}$ms reported in [33] yields the orange line.
@@
We highlight that, unlike other charge enhancement techniques [43, 45, 44], the rf electron cascade retains the non-demolition nature of in-situ dispersive readout measurements since the DQD system remains in an eigenstate after a measurement is performed [53, 1]. We note that the rf excitation is continuously applied for all measurements in this Article, we discuss the potential impact of this in the echo sequence section.
@@
Having established a method to distinguish between singlet and triplet spin states, our goal is to prepare and coherently control spin states of the DQD through voltage pulses along the detuning axis, $\epsilon$. Such pulses bring the DQD: i) from the (0,2) charge configuration in which a singlet is prepared; ii) into the (1,1) region where the electron spins are spatially separated between QDs and may evolve; and iii) back to an intermediate point where they can be measured (see Fig. 3(a-c)). Deep in the (1,1) region, the spin basis states are predominantly $\ket{\uparrow\downarrow}$ and $\ket{\downarrow\uparrow}$. Under an adiabatic ramp to $\epsilon=0$ for readout, these two states map onto the |$\ket{\text{T}_{0}(1,1)}$ and |$\ket{\text{S}(0,2)}$, respectively. The basis states are separated in energy by $h\Omega=\sqrt{J(\varepsilon)^{2}+\Delta E_{z}^{2}}$, where we include the kinetic exchange interaction $J(\varepsilon)$ and the Zeeman energy difference between electrons in each dot $\Delta E_{z}$. The spin detuning $\Delta E_{z}=\Delta g\mu_{B}B+g\mu_{B}\Delta B_{\text{HF}}$ (where $\mu_{B}$ is the Bohr magneton, $h$ is Planck’s constant) contains two main contributions: (i) the difference in g-factor between QDs $\Delta g=\left|g_{2}-g_{1}\right|$ arising from variations in the spin-orbit interaction (SOI) present near the Si/SiO2 interface  [54, 55, 26]; and (ii) the difference in the effective 29Si nuclear magnetic field experienced by each QD, $\Delta B_{\text{HF}}$. The random fluctuations in the effective magnetic field experienced by each electron in the DQD can be described by a normal distribution with mean of 0 (given the negligible spin polarisation) and standard deviation $\sigma_{\text{HF}}=30\pm 4~{}\upmu$T, as we shall see later. This value corresponds to a hyperfine energy strength of $3.4\pm 0.4$ neV, which aligns well with other reports in natural silicon [22, 48, 56, 57].
@@
The entangling two-qubit gate achieved between the spin qubits under the exchange interaction depends on the ratio $J/\Delta E_{z}$, tending to a $\sqrt{\text{SWAP}}$ operation as $J\gg\Delta E_{z}$ or a C-PHASE when $J\ll\Delta E_{z}$, though any gate within this set parameterised by $J/\Delta E_{z}$ can be used as the building block for a quantum error correcting code such as the surface code [63]. Defining the qubit quality factor as $Q=T_{2}^{*}\Omega$ (i.e. the number of periods before the amplitude of oscillations decays by $1/e$), we find $Q\gtrapprox 10$ in the region $J/\Delta E_{z}=2.1-3.2$, on par with prior reports across a range of semiconductor platforms [26, 47, 48, 58, 56]. This provides an upper bound estimate on the achievable two-qubit gate fidelity using the approximation $\mathcal{F}\approx 1-1/4Q\lessapprox 98\%$ [64]. To implement error-correctable two-qubit gates this fidelity would need to surpass $99\%$ [15], which could be achieved using isotopically enriched silicon. In the next section, we extend the coherence time using spin refocusing techniques.
@@
In the exchange echo sequence, after initialising a |$\ket{\text{S}(1,1)}$ state and applying a $X^{\prime}_{\pi/2}$ rotation, the two-electron system dephases under the effect of charge noise for a time $\tau_{1}$. The free evolution occurs at a detuning point where $J/\Delta E_{z}=12.8\pm 1.6$ where we measured $T_{2}^{*}=43\pm 3~{}$ns (see Fig 4(f)). We then refocus the spins by applying a $X^{\prime}_{\pi}$ rotation and let the system evolve for $\tau_{2}$ until a second $X^{\prime}_{\pi/2}$ rotation maps the resulting state to the |$\ket{\text{S}}$-|$\ket{\text{T}_{0}}$ axis. We extract the amplitude of the echo by fitting the signal in the $\tau_{2}-\tau_{1}$ domain (see Fig. 5(c)), and plot its value as a function of total free evolution time $\tau_{1}+\tau_{2}$. Fig. 5(d) shows the echo amplitude decays exponentially with total time ($\tau_{2}+\tau_{1}$), yielding a characteristic $T_{2,J}^{\text{echo}}=0.42\pm 0.02~{}\upmu$s which corresponds to an order of magnitude increase in the coherence time, similar to prior work with silicon devices [26, 56]. From our fit to Eq. (3), we extract a magnetic-noise-limited $T_{2,\Delta E_{z}}^{*}=3.3~{}\upmu$s at this set point, indicating that sources other than magnetic noise limit $T_{2,J}^{\text{echo}}$. This finding is in agreement with previous reports, which show that $T_{2,J}^{\rm echo}$ is an order of magnitude shorter than the magnetic noise limit [26, 56] (see Supplementary Table 2). We speculate that residual high-frequency charge noise limits the echo coherence [7]. This noise may include factors such as the effect of the rf readout excitation, which is active throughout the control sequence. We estimate the upper-bound on detuning fluctuations associated with this rf excitation to be $\delta\varepsilon_{\rm rms}^{\rm rf}\leq 2.7~{}\upmu{\rm eV}$ (see Methods Eq. (8)). However, this limitation is not intrinsic to the the readout scheme itself, as the effect can be effectively mitigated by implementing a pulsed readout protocol. In this approach, the rf-drive is deactivated during the control sequence, and only activated during the readout phase [65].